# Nonstabilizerness of a Boundary Time Crystal

### v3

@@ II Model and unraveling protocols
$\partial_{t}\hat{\rho}=-i\omega_{0}\left[\hat{S}_{x},\hat{\rho}\right]+\frac{\kappa}{S}\left(\hat{S}_{-}\hat{\rho}\hat{S}_{+}-\frac{1}{2}\left\{\hat{S}_{-}\hat{S}_{+},\hat{\rho}\right\}\right),$ (2)
@@ II.1 Quantum Jumps
The quantum jump (QJ) unraveling corresponds to a photodetection measurement scheme, wherein the detector performs projective measurements of the bath in the photon number basis {|$\{\ket{n_{B}}\}$, with $\hat{a}^{\dagger}\hat{a}\,\ket{n_{B}}=n_{B}\ket{n_{B}}$ [93]. Over a small time interval $\delta t$, the system evolves as follows: If no photon is detected, the state evolves deterministically according to the non-Hermitian Hamiltonian
@@
$\delta\ket{\psi}$ $=\left(-i\hat{H}_{\text{nH}}\,\delta t+\frac{\delta p}{2}\right)\ket{\psi}$
$\quad+dN\left(\frac{\hat{L}}{\sqrt{\langle\psi|\hat{L}^{\dagger}\hat{L}|\psi\rangle}}-\hat{\mathbb{I}}\right)\ket{\psi},$ (3)
@@ II.2 Quantum State Diffusion
$\delta\ket{\psi}$ $=\Bigl{[}-i\hat{H}\,\delta t-\frac{1}{2}\Bigl{(}\hat{L}^{\dagger}\hat{L}+|\ell|^{2}-2\ell^{*}\hat{L}\Bigr{)}\,\delta t$
$\quad\quad+(\hat{L}-\ell)\,\delta W\Bigr{]}\ket{\psi},$ (4)
@@ III Stabilizer entropy
where $\hat{P}=\sigma_{1}^{\alpha_{1}}\otimes\dots\otimes\sigma_{N}^{\alpha_{N}}$ are all the Pauli strings that constitute the Pauli group $\mathcal{P}_{N}$ of $N$ qubits. One can also define the density of SRE $m_{2}=\mathcal{M}_{2}/N$. For pure states $\hat{\rho}=\ket{\psi}\bra{\psi}$ |, |$\ket{\psi}$ is a stabilizer state if and only if $\hat{\rho}$ has precisely $2^{N}$ nonzero components in the operator basis given by the Pauli strings, while all other components are zero [72]. As such, the SRE provides a measure of the entropy associated with the probability distribution $\Pi_{P}=\mathrm{Tr}(\hat{P}\hat{\rho})^{2}/2^{N}$, adjusted by the entropy of a stabilizer state. If a mixed state can be produced by partially tracing a stabilizer pure state, its SRE is zero. The SRE obeys the bound $0\leq\mathcal{M}_{2}<N\ln 2$, where $\mathcal{M}_{2}=0$ if and only if the state is a stabilizer.
@@ III.1 Unravelling-independence of stabilizer entropy with negligible correlations
Our first result is that, for systems where correlations are negligible, the ensemble-average of the SRE $\mathcal{M}_{2}$ is unraveling-independent. This result particularly applies to the model of Eq. (2), for large system sizes, as we are going to show in the following. Here we propose a theoretical argument supporting this statement for generic systems without correlations. In this subsection, we denote the ensemble averages over the quantum trajectories as $\mathbb{E}[\bullet]$, whereas for the expectation values of an observable $\hat{P}$ over a state |$\ket{\psi^{(k)}}$ of the ensemble we use the superscript $(k)$, writing $P^{(k)}$.
@@ III.2 Nonstabilizerness and permutational invariance
where $\hat{\rho}_{N/2}$ is the reduced density matrix after partial tracing the degrees of freedom of half the system from the state $\hat{\rho}=\ket{\psi}\bra{\psi}$ |.
@@
To illustrate this point, we show one example of dynamics in Fig. 1 for a system of $N=40$ spins, in the BTC phase [$\omega_{0}=2\kappa$ in Eq. (2)]. The Lindblad dynamics is unraveled in terms of pure-state quantum trajectories, and averages are computed over $M=2000$ realizations for all considered unravelings, specified by the value of the parameter $\mu$ in Eq. (5). In panel (a), where we plot the nonstabilizerness density, the curves relative to the different unravelings are very close to each other (we have numerically verified that the difference between curves decreases by increasing $N$), whereas in panel (b) we see that the unravelings have completely different average bipartite entanglement. As expected, the magnetization along the $z$ axis, defined as $m_{z}=\langle\hat{S}_{z}\rangle/S$, is the same among all unravelings, panel (c), as for any other observable. Similar results are valid for the full distributions along quantum trajectories, as we show in Appendix A.
@@ IV Nonstabilizerness in the mean-field limit
where the magnetization components obey Eqs. (14). We note that, in the thermodynamic limit, the density of SRE is always finite, thus the stabilizer entropy always scales extensively with $N$, except when $m_{x}^{4}+m_{y}^{4}+m_{z}^{4}=1$. This only occurs when one of the magnetization components is equal to ±$\pm 1$ while the rest are zero, i. e., when the state is a stabilizer and is fully polarized along one of the three coordinate axes. We remark that Eq. (17) is only valid because correlations are negligible.
@@
Figure 3: Density of nonstabilizerness of the fixed point [$\overline{m_{2}}$, Eq. (18)] and average density of nonstabilizerness along the closed orbits around the fixed point [$\langle\widetilde{m_{2}}\rangle$, Eq. (20)] in the mean-field limit ($\Omega=\omega_{0}/\kappa$). The black dotted line is the numerical fit of Eq. (22). The red dashed line indicates the saturation value, estimated using Eq. (IV).
The SRE density is zero also for $\Omega=0$, where the fixed point is a steady state and is the dark state $|\ket{\psi_{D}}=\ket{\downarrow\downarrow\cdots\downarrow}$ (i. e., a pure state such that $\hat{L}\ket{\psi_{D}}=0$, also a stabilizer state), and $\Omega\to\infty$, where fixed points are the eigenstates of $\hat{H}=\omega_{0}\hat{S}_{x}$. The maximum SRE density is found for $\Omega^{*}_{1}=1/\sqrt{2}$ and $\Omega^{*}_{2}=\sqrt{2}$ and is equal to $\overline{m_{2}}^{*}=2\log 2-\log 3$, which coincides with the stabilizer entropy of the $T$ state |$\ket{T}\sim\ket{\uparrow}+\exp(i\pi/4)\ket{\downarrow}$.
@@
[…]
@@ V Finite-size results
Figure 4: Density of nonstabilizerness of the average state evolving according to the Lindblad equation [Eq. (2)] as function of time, for several system sizes $N$, compared with the mean-field result. (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is the fully polarized state $|\ket{\psi(0)}=\ket{\uparrow\uparrow\dots\uparrow}$.
@@ V.1 Nonstabilizerness of the average state
Figure 5: Purity of the average state evolving according to the Lindblad equation [Eq. (2)] as function of time, for several system sizes $N$. (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is the fully polarized state $|\ket{\psi(0)}=\ket{\uparrow\uparrow\dots\uparrow}$.
@@ Appendix A Nonstabilizerness, entanglement and magnetization distributions across the unravellings
To gain further insights into the observation that the stabilizer entropy is unraveling-independent while the entanglement entropy is not, we also studied the stabilizer entropy distributions on the different ensembles of quantum trajectories shown in Fig. 1. For instance, in Fig. 9(a) we show the histogram of the density of SRE of the $M=2000$ trajectories of the three ensembles, at $\kappa t=8$ (marked in the panels of Fig. 1 with vertical dotted lines). The histograms represent the probability density of a certain value of the nonstabilizerness density $m_{2}^{(k)}$ on the state |$\ket{\psi_{k}}$ of the unraveling considered. Besides having a similar mean, the distributions look identical. This is in line with the behavior of the magnetization, see Fig. 9(c), and completely different from the entanglement entropy, Fig. 9(b). These observations equally apply to other times $\kappa t$.

### v4

@@ II Model and unraveling protocols
[…]
@@ II.1 Quantum Jumps
[…]
@@
[…]
@@ II.2 Quantum State Diffusion
[…]
@@ III Stabilizer entropy
[…]
@@ III.1 Unravelling-independence of stabilizer entropy with negligible correlations
[…]
@@ III.2 Nonstabilizerness and permutational invariance
where $\hat{\rho}_{N/2}$ is the reduced density matrix after partial tracing the degrees of freedom of half the system from the state $|\hat{\rho}=\ket{\psi}\bra{\psi}$.
@@
To illustrate this point, we show one example of dynamics in Fig. 1 for a system of $N=40$ spins, in the BTC phase $\omega_{0}=2\kappa$ in Eq. (2)]. The Lindblad dynamics is unraveled in terms of pure-state quantum trajectories, and averages are computed over $M=2000$ realizations for all considered unravelings, specified by the value of the parameter $\mu$ in Eq. (5). In panel (a), where we plot the nonstabilizerness density, the curves relative to the different unravelings are very close to each other (we have numerically verified that the difference between curves decreases by increasing $N$), whereas in panel (b) we see that the unravelings have completely different average bipartite entanglement. As expected, the magnetization along the $z$ axis, defined as $m_{z}=\langle\hat{S}_{z}\rangle/S$, is the same among all unravelings, panel (c), as for any other observable. Similar results are valid for the full distributions along quantum trajectories, as we show in Appendix A.
@@ IV Nonstabilizerness in the mean-field limit
where the magnetization components obey Eqs. (14). We note that, in the thermodynamic limit, the density of SRE is always finite, thus the stabilizer entropy always scales extensively with $N$, except when $m_{x}^{4}+m_{y}^{4}+m_{z}^{4}=1$. This only occurs when one of the magnetization components is equal to $\pm 1$ while the rest are zero, i. e., when the state is a stabilizer and is fully polarized along one of the three coordinate axes. We remark that Eq. (17) is only valid because correlations are negligible.
@@
Figure 3: Density of nonstabilizerness of the fixed point $\overline{m_{2}}$, Eq. (18)] and average density of nonstabilizerness along the closed orbits around the fixed point $\langle\widetilde{m_{2}}\rangle$, Eq. (20)] in the mean-field limit ($\Omega=\omega_{0}/\kappa$). The black dotted line is the numerical fit of Eq. (22). The red dashed line indicates the saturation value, estimated using Eq. (IV).
[…]
@@
$\langle\widetilde{m_{2}}\rangle^{(\Omega\to\infty)}$ $\approx-\frac{1}{4\pi}\int_{0}^{\pi}\sin\theta d\theta\int_{0}^{2\pi}d\phi$
@@ V Finite-size results
Figure 4: Density of nonstabilizerness of the average state evolving according to the Lindblad equation [Eq. (2)] as function of time, for several system sizes NN, compared with the mean-field result. (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is the fully polarized state $|\ket{\psi(0)}=\ket{\uparrow\uparrow\dots\uparrow}$.
@@ V.1 Nonstabilizerness of the average state
Figure 5: Purity of the average state evolving according to the Lindblad equation [Eq. (2)] as function of time, for several system sizes NN. (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is the fully polarized state $|\ket{\psi(0)}=\ket{\uparrow\uparrow\dots\uparrow}$.
@@ Appendix A Nonstabilizerness, entanglement and magnetization distributions across the unravellings
[…]